OpenStudy (anonymous):

Find the y-intercept of x – 5y = 10. A. (-2, 0) B. (0, -2) C. (0, 10) D. (10, 0)

OpenStudy (amistre64):

yint mean x=0, so that eliminates 2 options

OpenStudy (amistre64):

when x=0; what does y= for the statement to be true?

OpenStudy (anonymous):

2

OpenStudy (amistre64):

close, and accepting for a typo; yes 0 -5y = 10 -5y = 10 ; when y= -2

OpenStudy (amistre64):

we know it cant be 10 regardless, so yeah; B is a safe bet

OpenStudy (anonymous):

Thats what I though I just wanted to make shore I am right

OpenStudy (amistre64):

math ability has to start somewhere :) i still miss signs and confuse adding for multiplication :/